# CYCLO: Cyclic Graph Transformer Approach to Multi-Object Relationship Modeling in Aerial Videos

### v1

@@ 1 Introduction
Visual scene understanding has shown significant progress in extracting semantic information from images and videos using deep learning algorithms [1, 2]. Building upon the significant progress in visual scene understanding using deep learning algorithms [1, 2], video scene graph generation (VidSGG) extends the concept of Scene Graph Generation (SGG) from static images to dynamic video, representing object relationships within a graph structure that evolves over time. VidSGG [3, 4, 5] focus on the temporal dimension by constructing a dynamic graph structure, encapsulating the spatial and temporal relationships among object interactions across frames. This helps in understanding human-object interactions [6, 7], temporal events [8, 9, 10], and reasoning [11, 12]. However, drone-captured videos present unique challenges due to larger image sizes and higher object density in Unmanned Aerial Vehicle (UAV) datasets [13, 14, 15]. Despite recent advances in tiny object detection [16, 17, 18], current algorithms still need to effectively model object interactions and their temporal evolution in aerial videos, which have various applications in surveillance, disaster response
@@
In this paper, we introduce AeroEye, the first dataset for Video Scene Graph Generation in drone-captured videos featuring Aerial-Oblique-Eye views. AeroEye distinguishes itself by showcasing a rich tapestry of aerial videos and an extensive set of predicates describing the intricate relations and positions of multi-objects. To address multi-object relationship modeling in aerial videos from the AeroEye dataset, we propose the Cyclic Graph Transformer (CYCLO). This new approach can establish circular connectivity among frames and enables the model to capture direct and long-range temporal relationships. By continuously updating history across a ring topology, CYCLO allows the model to handle sequences with inherent cyclic patterns, facilitating the processing of object relationships in the correct temporal order. Furthermore, CYCLO provides several advantages to VidSGG, including the ability to model periodic and overlapping relationships, predict object interactions by reasoning from previous cycles, facilitate information transfer across frames, and efficiently utilize long sequences, addressing the limitations of previous methods [19, 5]. They usually struggle with long-term dependencies due to the diminishing influence of inputs over time.
@@
Contributions of this Work. There are three main contributions to this work. First, we introduce a new AeroEye dataset for VidSGG in drone videos, augmented with numerous predicates and diverse scenes to capture the complex relationships in aerial videos. Second, we propose the CYCLO approach, utilizing circular connectivity among frames to enable periodic and overlapping relationships. It allows the model to capture long-range dependencies and process object interactions in the appropriate temporal arrangement. Finally, the proposed CYCLO approach outperforms prior methods on two large-scale in-the-wild VidSGG datasets, including PVSG [4] and ASPIRe [5]. Interestingly, using the same method (e.g., our CYCLO), the ratio of correct predictions to incorrect predictions (R/mR) on AeroEye is higher than PVSG (Tables 4 and  6), despite having more predicates (Table A.11) and tiny objects. This suggests that our dataset is less visually ambiguous than PVSG.
@@ 2 Related Work
Table 1: Comparison of available datasets on drone images and videos. Viewpoints include drone-captured perspectives, which is our main focus in this work, including aerial (top-down), oblique (slanted), and ground (eye-level) perspectives. # denotes the number of corresponding items. The best values are highlighted. Full comparison with other scene graph generation benchmarks is in Table A.11
@@
BBox Relation #Instances aerial oblique ground
DOTA [13] - 11.3K 1490 $\times$ 957 2.2M 18 - -
@@
AeroEye (Ours) 2.3K 261.5K 3840 $\times$ 2160 2.2M 56 384 29
CC-BY License of Creative Commons
In this section, we first review the prior Visual Scene Graph Generation datasets and benchmarks. Then, we summarize the related Video Scene Graph Generation problems.
@@ 2.1 Visual Scene Graph Generation Datasets and Benchmarks
Benchmarks. The existing benchmark focuses on Image Scene Graph Generation (ImgSGG) and Video Scene Graph Generation (VidSGG). ImgSGG identifies and categorizes relationships between objects within an image into predefined relational categories, including Transformer-based methods [37, 38, 39, 40] and generative-based models [41, 42, 43]. VidSGG leverages the dynamic nature of object interactions over time to better identify relationships, as the temporal dimension of videos provides a richer context for understanding semantic connections within the scene. Current methods using hierarchical structures [44, 5] or Transformer architectures [19, 45, 46] excel at capturing long-range dependencies and complex interactions, advancing video understanding in video captioning[8, 9, 10], visual question answering [11, 12], and video grounding [47, 48, 49].
@@ 2.2 Video Scene Graph Generation
VidSGG can be categorized into two main types based on the granularity of its graph representation. Video-level SGG represents object trajectories as graph nodes, capturing constant relations between objects for a video. Various methods have been proposed to address this problem, incorporating Conditional Random Fields [50], abstracting videos [51], and iterative relation inference techniques [52] on fully connected spatio-temporal. However, focusing primarily on recognizing video-level relations directly based on object-tracking results and neglecting frame-level scene graphs results in a cumbersome pipeline highly dependent on tracking accuracy. In contrast, Frame-level SGG defines the graph at the frame level, allowing relations to change over time. The releases of the benchmark datasets [53, 4, 5] have prompted the development of VidSGG models. TRACE [44], for instance, employs a hierarchical relation tree to capture spatio-temporal context information, while CSTTran [19] uses a spatio-temporal transformer to solve the problem. Recently, hierarchical interlacement graph (HIG) [5] abstracts relationship evolution using a sequence of hierarchical graphs.
@@ 2.3 Discussions
Limitations. While the batch-progressive approach considers temporal information, both these progressive and batch-progressive approaches have limitations in modeling the full complexity of temporal dynamics and dependencies in the video. In contrast, the hierarchical graph approach can capture complex interactions and relationships between objects by considering the temporal evolution of graphs at different granularity levels. However, the hierarchical graph requires observation of the entire video to construct the graph (i.e., offline method).
@@ 3.1 Dataset Collection
Relationship Classes and Instance Formulation. In Fig. 3, we focus on two aspects of object relationships: positions (e.g. in front of, behind, next to) and relations, which consist of movement actions (e.g. chasing, towing, overtaking) and collision actions (e.g. hitting, crashing, colliding). These relationships are semantically complex and require detailed spatio-temporal context reasoning for recognition. Following previous datasets [3, 4, 5], we define relationship instances at the frame level, considering the long-term spatial-temporal context. Each instance is formulated as a triplet $\textlangle s,o,p\textrangle$, where $s$ and $o$ denote the bounding boxes of the subject and object, and $p$ represents the predicate, included in Table A.8 and Table A.9 and summarized in Fig. 4.
@@ 3.2 Data Specification
Data Annotation. We annotate keyframes at 5FPS to capture frequent and rapid changes in positions and relations in aerial videos, reducing redundancy while keeping up with interaction changes. Our two-stage annotation pipeline first performs object localization and tracking and then relationship instance annotation. Although we annotate relationship instances frame by frame illustrated in Fig. 3, we easily create relationship tubes using the provided object tracking ID by connecting the same pair of objects with the same relationship predicate across consecutive frames. The annotation file includes object information (i.e., bounding boxes, category names, and tracking IDs) and relationships within each frame. The quality control and annotation examples are discussed in Section A.2 of Appendices.
Data Statistics. The AeroEye dataset is a collection of 2,260 videos with 261,503 frames, annotated with over 2.2 million bounding boxes across 56 object categories typically observed from aerial, oblique, and ground perspectives captured by drone. Specifically, our AeroEye dataset consists of 384 predicates divided into two relationship categories: 135 positions and 249 relations. The key strength of AeroEye is its richness in relationships. On average, each video in the dataset has 127 frames, providing moderate temporal depth for capturing detailed interactions. The average number of frames per scene is 8,970, indicating substantial variability and complexity. AeroEye is rich in relationships, with more than 43 million relationship instances. In Table 1, we provide a detailed comparison with related datasets, while Fig. 5 presents statistics on the AeroEye dataset. In addition, predicate definitions and additional statistics are disccused in Sections A.1 and A.4 of the Appendices.
@@ 4 The Proposed Approach
In this section, we introduce the CYCLO approach, which consists of two main components, including the Spatial Attention Graph and the Cyclic Temporal Graph Transformer.
@@ 4.2 Spatial Attention Graph
Self-attention mechanisms in one-stage object detectors [56, 16] model relationships between objects, allowing insights into the dynamics between entities without relying on additional contextual information. For example, in an aerial parking lot video with cars, vans, and people, if the self-attention layer strongly connects the queries representing the person and the car, it suggests an interaction, such as the person entering the vehicle. Inspired by [40], in our CYCLO approach, to construct the static graph in each frame $t$, we utilize the DETR decoder to establish bidirectional connections among object queries. In particular, we compute relational representations at each layer $l$ by concatenating the query and key vectors, $Q_{t}^{l}$ and $K_{t}^{l}$, for every object pair. This process ensures the layer $l-1$ output seamlessly transitions as input to layer $l$. We omit the superscript $L$ related to the final layer to simplify the following discussion. At each frame, $\widehat{R}_{a,t}^{l}$ captures the dynamic interplay of relations at layer $l$, utilizing their query and key vectors. Furthermore, $\widehat{R}_{z,t}$ leverages object queries in the final layer for object detection. These relationships are formally defined as in Eqn. (1).
@@ 4.3 Cyclic Temporal Graph Transformer
In Eqn. (3), $\eta$ is a shift term enabling cyclical indexing via ($\mathrm{mod}$ $T$). This cyclical indexing, illustrated in Fig. 6, allows for continuous sequence processing by connecting the end to the beginning, which is crucial for predicting movements in dynamic interactions where past events influence future actions (e.g. a car navigating a roundabout). Different from the standard self-attention (Eqn. (B.21)), which is permutation equivariant without positional encodings, the cyclical indexing in cyclic attention that is non-permutation equivariant. This property is crucial for multi-object relationship modeling as discussed in Subsection 4.2. Detailed explanations are included in Section B.4 of the Appendices.
Temporal Graph Transformer. Our Temporal Graph Transformer refines spatial attention graphs, $\{\widehat{G}_{t}\}_{t=1}^{T}$, into a sequence of dynamic graphs, $\{G_{t}\}_{t=1}^{T}$, leveraging the temporal dynamics and spatial interactions of objects across video frames. Our approach employs a series of cyclic attention blocks configured within multi-head attention to refine object representations by integrating features from adjacent frames. The core of our approach is the integration of cyclic attention into a multi-head structure, which processes the sequence of input features $\widehat{Z}=\{\widehat{Z}_{t}\}_{t=1}^{T}$, represented as in Eqn. (4).
@@
Loss Function. Visual object relationships often involve similar predicates (e.g. parking next to and stopping next to). Thereby, we use multi-label margin loss, as in [19]:
$\small\mathcal{L}_{p}(r,\mathcal{P}^{+},\mathcal{P}^{-})=\sum_{p\in\mathcal{P}^{+}}\sum_{q\in\mathcal{P}^{-}}\max(0,1-\phi(r,p)+\phi(r,q))$ (5)
Here, $r$ is a subject-object pair, with $\mathcal{P}^{+}$ and $\mathcal{P}^{-}$ representing positive and negative predicates, respectively. Additionally, object distributions are determined using neural networks with ReLU and batch normalization, applying cross-entropy loss. The total loss combines these two losses.
@@ 5.1 Implementation Details
Settings. We employ DINO [16] to extract the spatial attention graphs (in Section 4.2). DINO is trained with ResNet-50 backbone and $1500$ queries on MAVREC, achieving $92.35$ mAP on the validation set. The pre-trained detector is applied to baselines, and parameters are fixed during subsequent task training. Our model is trained on 8 $\times$ A600 GPUs using 12 epochs with AdamW optimizer (initial learning rate of $1e^{-5}$ and a batch size of $1$), gradient clipping (max norm of $5$).
Evaluation Metrics. We evaluate models on two standard tasks in image-based scene graph generation followed by previous work [59, 4] that are predicate classification (PredCls), scene graph classification (SGCls), and scene graph detection (SGDet). While SGCls predicts relationships given ground truth objects, SGDet involves detecting objects and predicting relationships. These tasks are evaluated using Recall (R@K) and mean Recall (mR@$K$), where $K\in\{10,20,50\}$.
5.2 Ablation Study
Table 2: Our performance (%) on AeroEye for varying shift terms at Recall (R) and mean Recall (mR).
@@ 5.2 Ablation Study
Semantic Dynamics in Cyclic Attention. By altering $\eta$ (in Eqn. (3)), we consider the permutation or non-equivariance equivariance. If the predictions systematically adapt to the shifts induced by different $\eta$ values, it demonstrates a degree of permutation equivariance. Conversely, if the predictions change in ways that do not correspond systematically to these shifts, it may indicate non-permutation equivariance. Table 2 shows a decrease performance, implying disrupted temporal patterns and non-permutation invariance, indicated by unpredictable output changes relative to input shifts.
@@
Cyclic Dependency. To further validate the cyclic dependency of our model, we discarded frames from every successive frame. Removing frames disrupts the temporal continuity of the sequence, which is crucial for maintaining the the cyclic nature of video. If the cyclic model presupposes that each frame has direct relationships with its adjacent frames circularly, then removing frames could sever these relationships, potentially diminishing the ability to leverage cyclical patterns effectively. Indeed, Table  3 reveals a decrease at Recall and mean Recall when frames were reduced.

### v4

# CYCLO [Uncaptioned image]: Cyclic Graph Transformer Approach to Multi-Object Relationship Modeling in Aerial Videos

@@ 1 Introduction
Visual scene understanding has shown significant progress in extracting semantic information from images and videos using deep learning algorithms [1, 2, 3]. Building upon the significant progress in visual scene understanding using deep learning algorithms [4, 5], video scene graph generation (VidSGG) extends the concept of Scene Graph Generation (SGG) from static images to dynamic video, representing object relationships within a graph structure that evolves over time. VidSGG [6, 7, 8] focus on the temporal dimension by constructing a dynamic graph structure, encapsulating the spatial and temporal relationships among object interactions across frames. This helps in understanding human-object interactions [9, 10], temporal events [11, 12, 13], and reasoning [14, 15]. However, drone-captured videos present unique challenges due to larger image sizes and higher object density in Unmanned Aerial Vehicle (UAV) datasets [16, 17, 18]. Despite recent advances in tiny object detection [19, 20, 21], current algorithms still need to effectively model object interactions and their temporal evolution in aerial videos, which have various applications in surveillance, disaster response.
@@
In this paper, we introduce AeroEye, the first dataset for Video Scene Graph Generation in drone-captured videos featuring Aerial-Oblique-Eye views. AeroEye distinguishes itself by showcasing a rich tapestry of aerial videos and an extensive set of predicates describing the intricate relations and positions of multi-objects. To address multi-object relationship modeling in aerial videos from the AeroEye dataset, we propose the Cyclic Graph Transformer (CYCLO). This new approach can establish circular connectivity among frames and enables the model to capture direct and long-range temporal relationships. By continuously updating history across a ring topology, CYCLO allows the model to handle sequences with inherent cyclic patterns, facilitating the processing of object relationships in the correct temporal order. Furthermore, CYCLO provides several advantages to VidSGG, including the ability to model periodic and overlapping relationships, predict object interactions by reasoning from previous cycles, facilitate information transfer across frames, and efficiently utilize long sequences, addressing the limitations of prior methods [22, 8]. They usually struggle with long-term dependencies due to the diminishing influence of inputs over time.
@@
Contributions of this Work. There are three main contributions to this work. First, we introduce a new AeroEye dataset for VidSGG in drone videos, augmented with numerous predicates and diverse scenes to capture the complex relationships in aerial videos. Second, we propose the CYCLO approach, utilizing circular connectivity among frames to enable periodic and overlapping relationships. It allows the model to capture long-range dependencies and process object interactions in the appropriate temporal arrangement. Finally, the proposed CYCLO approach outperforms prior methods on two large-scale in-the-wild VidSGG datasets, including PVSG [7] and ASPIRe [8]. Interestingly, using the same method (e.g., our CYCLO), the ratio of correct predictions to incorrect predictions (R/mR) on AeroEye is higher than PVSG (Tables 4 and  7), despite having more predicates (Table 1) and tiny objects. This suggests that our dataset is less visually ambiguous than PVSG.
@@ 2 Related Work
Table 1: Comparison of available datasets for scene graph generation. The top two blocks present image and video scene graph datasets, while the next two focus on image and drone video datasets. Viewpoints include five types: from ego (1st-person) view to 3rd-person view, and drone-captured perspectives, which is our main focus in this work (/ in colors), including aerial (top-down), oblique (slanted), and ground (eye-level) perspectives. # denotes the number of corresponding items. The best values in drone blocks are highlighted.
@@
BBox Relation #Instances 3rd ego aerial oblique ground
Visual Genome [23] - 108K - 3.8M 33K 42K -
VG-150 [24] - 88K - 2.8M 150 50 -
VrR-VG [25] - 59K - 282.4K 1.6K 117 -
GQA [26] - 85K - 1.7K 310 -
PSG [27] - 49K - 538.2K 80 56 -
VidVRD[28] 1K 296.2K 1920 $\times$ 1080 15.1K 35 132 -
Action Genome [6] 10K 234.3K 1280 $\times$ 720 476.3K 25 25 -
ASPIRe [8] 1.5K 1.6M 1280 $\times$ 720 167.8K 833 4.5K 7
SportsHHI [29] 80 11.4K 1280 $\times$ 720 118.1K 1 34 2
VidOR [30] 10K 55.4K 640 $\times$ 360 50K 80 50 1
VidSTG [31] 10K 55.4K 640 $\times$ 360 50K 80 50 1
EPIC-KITCHENS [32] 700 11.5K 1920 $\times$ 1080 454.3K 21 13 1
PVSG [7] 400 153K 1920 $\times$ 1080 7.6K 126 57 7
DOTA [16] - 11.3K 1490 $\times$ 957 1.8M 18 - -
@@
AeroEye (Ours) 2.3K 261.5K 3840 $\times$ 2160 1.8M 57 384 29
In this section, we review the existing datasets and benchmarks for Visual Scene Graph Generation, followed by a summary of the key challenges and issues related to Video Scene Graph Generation.
@@ 2.1 Visual Scene Graph Generation Datasets and Benchmarks
Benchmarks. The existing benchmark focuses on Image Scene Graph Generation (ImgSGG) and Video Scene Graph Generation (VidSGG). ImgSGG identifies and categorizes relationships between objects within an image into predefined relational categories, including Transformer-based methods [27, 48, 49, 50] and generative-based models [51, 52, 53]. VidSGG leverages the dynamic nature of object interactions over time to better identify relationships, as the temporal dimension of videos provides a richer context for understanding semantic connections within the scene. Current methods using hierarchical structures [54, 8] or Transformer architectures [22, 55, 56, 57] excel at capturing long-range dependencies and complex interactions, advancing video understanding in video captioning[11, 12, 13], visual question answering [14, 15], and video grounding [58, 59, 60].
@@ 2.2 Video Scene Graph Generation
VidSGG can be categorized into two main types based on the granularity of its graph representation. Video-level SGG represents object trajectories as graph nodes, capturing constant relations between objects for a video. Various methods have been proposed to address this problem, incorporating Conditional Random Fields [61], abstracting videos [62], and iterative relation inference techniques [63] on fully connected spatio-temporal. However, focusing primarily on recognizing video-level relations directly based on object-tracking [64, 65, 66] results and neglecting frame-level scene graphs results in a cumbersome pipeline highly dependent on tracking accuracy. In contrast, Frame-level SGG defines the graph at the frame level, allowing relations to change over time. The releases of the benchmark datasets [67, 7, 8] have prompted the development of VidSGG models. TRACE [54], for instance, employs a hierarchical relation tree to capture spatio-temporal context information, while CSTTran [22] uses a spatio-temporal transformer to solve the problem. Recently, hierarchical interlacement graph (HIG) [8] abstracts relationship evolution using a sequence of hierarchical graphs.
@@ 2.3 Discussions
Limitations. While the batch-progressive approach considers temporal information, both these progressive and batch-progressive approaches have limitations in modeling the full complexity of temporal dynamics and dependencies in the video. In contrast, the hierarchical graph approach can capture complex interactions and relationships between objects by considering the temporal evolution of graphs at different granularity levels. However, the hierarchical graph requires analyzing the entire video before constructing the graph (i.e. offline method). These limitations underscore the need for more advanced approaches to efficiently model temporal dynamics, adaptively update memory to handle evolving video data, and accurately capture the intricate relationships between objects.
@@ 3.1 Dataset Collection
Relationship Classes and Instance Formulation. In Fig. 3, we focus on two aspects of object relationships: positions (e.g. in front of, behind, next to) and relations, which consist of movement actions (e.g. chasing, towing, overtaking) and collision actions (e.g. hitting, crashing, colliding). These relationships are semantically complex and require detailed spatio-temporal context reasoning for recognition. Following previous PVSG benchmark datasets [6, 7, 8], we define relationship instances at the frame level, considering the long-term spatial-temporal context. Each instance is formulated as a triplet $\textlangle s,o,p\textrangle$, where $s$ and $o$ denote the bounding boxes of the subject and object, and $p$ represents the predicate (i.e. position and relation), included in Tables A.8A.9 which are summarized in Fig. 4. In addition, Fig. A.11 presents selected samples from our AeroEye dataset.
@@ 3.2 Data Specification
Data Annotation. We annotate keyframes at 5FPS to capture frequent and rapid changes in positions and relations in aerial videos, reducing redundancy while keeping up with interaction changes. Our two-stage annotation pipeline first performs object localization and tracking and then relationship instance annotation. To generate diverse predicates, we leverage the GPT4RoI [70] model, which combines visual and linguistic data to generate detailed descriptions of object relationships within specified regions of interest. Although we annotate relationship instances frame by frame, as illustrated in Fig. 3, we easily create relationship tubes using the provided object tracking ID by connecting the same pair of objects with the same relationship predicate across consecutive frames. The annotation file includes object information (i.e., bounding boxes, category names, and tracking IDs) and relationships within each frame. Details on the quality control process and annotation examples can be found in Section A.2 of the Appendices.
Data Statistics. The AeroEye dataset is a collection of 2,260 videos with 261,503 frames, annotated with over 2.2 million bounding boxes across 56 object categories typically observed from aerial, oblique, and ground perspectives captured by drone. Specifically, our AeroEye dataset consists of 384 predicates divided into two relationship categories: 135 positions and 249 relations. The key strength of AeroEye is its richness in relationships. On average, each video in the dataset has 127 frames, providing moderate temporal depth for capturing detailed interactions. The average number of frames per scene is 8,970, indicating substantial variability and complexity. AeroEye is rich in relationships, with more than 43 million relationship instances. In Table 1, we provide a detailed comparison with related datasets, while Fig. 5 presents statistics on the AeroEye dataset. In addition, predicate definitions and additional statistics are discussed in Sections A.1 and A.4 of the Appendices.
@@ 4 The Proposed Approach
In this section, we present our CYCLO approach, including the Spatial Attention Graph and the Cyclic Temporal Graph Transformer. The Spatial Attention Graph captures spatial dependencies and interactions between objects within the frame. In contrast, the Cyclic Temporal Graph Transformer models temporal relationships across frames, capturing short-term and long-term dynamics.
@@ 4.2 Spatial Attention Graph
Self-attention mechanisms in one-stage object detectors [71, 19] model relationships between objects, allowing insights into the dynamics between entities without relying on additional contextual information. For example, in an aerial parking lot video with cars, vans, and people, if the self-attention layer strongly connects the queries representing the person and the car, it suggests an interaction, such as the person entering the vehicle. Inspired by [50], in our CYCLO approach, to construct the static graph in each frame $t$, we utilize the DETR decoder to establish bidirectional connections among object queries. In particular, we compute relational representations at each layer $l$ by concatenating the query and key vectors, $Q_{t}^{l}$ and $K_{t}^{l}$, for every object pair. This process ensures the layer $l-1$ output seamlessly transitions as input to layer $l$. We omit the superscript $L$ related to the final layer to simplify the following discussion. At each frame, $\widehat{R}_{a,t}^{l}$ captures the dynamic interplay of relations at layer $l$, utilizing their query and key vectors. Furthermore, $\widehat{R}_{z,t}$ leverages object queries in the final layer for object detection. These relationships are formally defined in Eqn. (1).
@@ 4.3 Cyclic Temporal Graph Transformer
In Eqn. (3), $\eta$ is a shift term enabling cyclical indexing via $\mathrm{mod}$ $T$. The cyclical indexing, illustrated in Fig. 6, allows for continuous sequence processing by connecting the end to the beginning, which is crucial for predicting movements in dynamic interactions where past events influence future actions (e.g. a car navigating a roundabout). Cyclical indexing differs from standard self-attention as it is a permutation equivariant without positional encodings. In contrast, cyclical attention is non-permutation equivariant, which depends on the original sequence order. This property is crucial for multi-object relationship modeling, where maintaining the chronological order of interactions is essential. For instance, in a surveillance scenario, the sequence of a car stopping for a pedestrian must preserve the order of events, ensuring the vehicle stops before the pedestrian appears.
Temporal Graph Transformer. Our Temporal Graph Transformer refines spatial attention graphs, $\{\widehat{G}_{t}\}_{t=1}^{T}$, into a sequence of dynamic graphs $\{\mathcal{G}_{t}\}_{t=1}^{T}$, leveraging the temporal dynamics and spatial interactions of objects across video frames. Our approach employs a series of cyclic attention blocks configured within multi-head attention to refine object representations by integrating features from adjacent frames. The core of our approach is the integration of cyclic attention into a multi-head structure, which processes the sequence of input features $\widehat{Z}=\{\widehat{Z}_{t}\}_{t=1}^{T}$, represented as in Eqn. (4).
@@
Loss Function. Visual object relationships involve predicates that may appear quite similar, such as “parking next to” and “stopping next to”. Thereby, we utilize a multi-label margin loss, as in [22]:
$\mathcal{L}_{p}(r,\mathcal{P}^{+},\mathcal{P}^{-})=\sum_{p\in\mathcal{P}^{+}}\sum_{q\in\mathcal{P}^{-}}\max(0,1-\phi(r,p)+\phi(r,q))$ (5)
In Eqn. (5), $r$ represents a subject-object pair, and $\mathcal{P}^{+}$ and $\mathcal{P}^{-}$ correspond to positive and negative predicates, respectively. The term $\phi(r,p)$ measures the compatibility of the pair $r$ with the predicate $p$. Additionally, object distributions are modeled using neural networks with ReLU activation and batch normalization. Cross-entropy loss is applied during the learning process. The total loss is a combination of the multi-label margin loss $\mathcal{L}_{p}$ and the cross-entropy loss $\mathcal{L}_{ce}$, defined as:
$\mathcal{L}_{total}=\mathcal{L}_{p}+\lambda\mathcal{L}_{ce}$ (6)
where $\lambda$ is a weight balancing the contribution of the cross-entropy loss $\mathcal{L}_{ce}$.
@@ 5.1 Implementation Details
Table 2: Our performance (%) on AeroEye with shift values ($\eta$ in Eqn. (3)) at Recall (R) and mean Recall (mR).
@@
Settings. We employ DINO [19] to extract the spatial attention graphs (in Section 4.2). DINO is trained with ResNet-50 backbone and $1500$ queries on MAVREC, achieving $92.35$ mAP on the validation set. The pre-trained detector is applied to baselines, and parameters are fixed during subsequent task training. Our model is trained on 8 $\times$ A6000 GPUs using 12 epochs with AdamW optimizer (initial learning rate of $1e^{-5}$ and a batch size of $1$), gradient clipping (max norm of $5$).
Evaluation Metrics. We evaluate models on two standard tasks in image-based scene graph generation followed by previous work [74, 7] that are predicate classification (PredCls), scene graph classification (SGCls), and scene graph detection (SGDet). While SGCls predicts relationships given ground truth objects, SGDet involves detecting objects and predicting relationships. These tasks are evaluated using Recall (R@K) and mean Recall (mR@$K$), where $K\in\{20,50,100\}$.
5.2 Ablation Study
Semantic Dynamics in Cyclic Attention. By altering $\eta$ (in Eqn. (3)), we consider the permutation or non-equivariance equivariance. If the predictions systematically adapt to the shifts induced by different $\eta$ values, it demonstrates a degree of permutation equivariance. Conversely, if the predictions change in ways that do not correspond systematically to these shifts, it may indicate non-permutation equivariance. Table 2 shows a decrease in performance, implying disrupted temporal patterns and non-permutation invariance, indicated by unpredictable output changes relative to input shifts.
Cyclic Dependency. To further validate the cyclic dependency of our model, we discarded frames from every successive frame. Removing frames disrupts the temporal continuity of the sequence, which is crucial for maintaining the the cyclic nature of video. If the cyclic model presupposes that each frame has direct relationships with its adjacent frames circularly, then removing frames could sever these relationships, potentially diminishing the ability to leverage cyclical patterns effectively. Indeed, Table  3 informs a decrease in Recall and mean Recall when frames were reduced.